OpenStudy (anonymous):

Shawn makes $235 a day on a job. How much will he earn for working 1/2 day?

OpenStudy (anonymous):

($235 per day) * (1/2) days = $117.5

OpenStudy (anonymous):

A garden measures 6 ft by 7 ft needs fencing. If the fencing costs $8 per foot, how much will it cost to put fencing around the garden?

OpenStudy (anonymous):

you need four fences to go around the garden. one is 6 ft, the next is 7 ft, the next is 6 ft, the next is 7ft. 6+7+6+7 = 26 ft. $8 p/foot * 26ft = $8*26 =$208